OpenStudy (anonymous):

kindly help me integrate this please. zdz/(z^2+4z+4) i think i have a partial fraction here.

OpenStudy (unklerhaukus):

\[\int \frac{z\cdot\text dz}{z^2+4z+4}\]

OpenStudy (unklerhaukus):

\[\frac{z}{z^2+4z+4}=\frac z{(z+2)^2}=\frac A{z+2}+\frac B{(z+2)^2}\]

OpenStudy (anonymous):

{z/(z+2)(z+2)} {z/(z+2)^2} 1/z+2 - 2/(z+2)^2 integrating ln(z+2) +2/z+2 i think that i have done it right

OpenStudy (unklerhaukus):

\[z=(z+2)A+B\] \[z=0\Rightarrow0=2A+B\] \[\qquad\qquad\qquad z=-2 \Rightarrow -2=B\] \[0=2A-2\rightarrow 1=A\] \[\frac{z}{z^2+4z+4}=\frac z{(z+2)^2}=\frac 1{z+2}-\frac 2{(z+2)^2}\] \[\int\frac1{z+2}-\frac 2{(z+2)^2}\text dz\]

OpenStudy (unklerhaukus):

\[\int\frac1{z+2}-\frac 2{(z+2)^2}\text dz\] \[=\ln(z+2)+\frac2{z+2}+c\]

OpenStudy (amistre64):

\[\int\frac{z}{z^2+4z+4}dz\] \[\int\frac{z+(2-2)}{z^2+4z+4}dz\] \[\int\frac{2}{2}*\frac{z+2-2}{z^2+4z+4}dz\] \[\frac{1}{2}\int\frac{2z+4-4}{z^2+4z+4}dz\] \[\frac{1}{2}\int\frac{2z+4}{z^2+4z+4}dz-\frac42\int\frac{1}{(z+2)^2}dz\] \[ln(z+2)+2 (z+2)^{-1}+C\]